# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ VI-D Personalized Education
Personalized education considers individual learning styles, preferences, and pace to enhance student engagement, foster a deeper interest in learning, and ultimately improve academic performance. The integration of wireless networks and LLMs has the potential to revolutionize personalized education by providing tailored learning experiences for students. Through real-time data transmission from devices, educators can adjust teaching materials based on the progress of different students, addressing diverse learning needs. For example, personalized learning platforms can deliver content suited to various learning styles, with visual learners receiving more graphics and video materials, while auditory learners engage with audio lectures. Teachers can analyze students’ learning data in real time to identify those needing additional support and provide targeted guidance. In addition, adaptive learning systems can dynamically modify learning paths according to students’ performance, ensuring they learn at an appropriate difficulty level. Furthermore, educational institutions can develop commercial personalized education tools to optimize curriculum settings and resource allocation, promoting the widespread adoption and advancement of personalized education. By leveraging these strategies, the combination of wireless networks and LLM would significantly enhance the effectiveness of personalized education, allowing every student to benefit from a customized learning experience.
VII Future Directions
@@
In this section, we present a vision for the future directions of the development of large language models (LLMs) for wireless networks, including quantum LLMs, on-device LLMs, neural-symbolic LLMs, and autonomous AI agents.
VII-A Quantum LLMs
@@
Quantum computing [214] is a new computing paradigm that utilizes the principles of quantum mechanics to manipulate quantum information units for computation. By employing quantum bits (qubits) to process information in parallel, it achieves greater computational power and the capacity to solve more complex problems than traditional computing methods. The advancement of quantum computing is anticipated to have a profound impact on existing technologies, including LLMs. For instance, quantum computing has the potential to significantly improve the efficiency of computing, improve data security, and speed up model training, thereby greatly boosting the performance of LLMs. However, to fully realize the benefits of quantum computing, it is essential to transform conventional computer algorithms into quantum algorithms specifically designed for quantum architectures. Research on quantum-enhanced LLMs needs to prioritize optimizing model architecture design, model training and inference, and resource allocation to maximize the advantages provided by quantum computing. In addition, exploring quantum-inspired neural networks and quantum machine learning frameworks could unlock new paradigms in language understanding, reasoning, and data processing.
VII-B On-Device LLMs
@@
LLMs must evolve beyond simple response generation to deeply integrate into machine learning (ML) workflows while operating efficiently on edge devices. They should assist with complex ML-specific tasks such as model selection, hyperparameter tuning, feature engineering, and AutoML while leveraging multimodal capabilities to process diverse data types, including text, images, and structured datasets. However, deploying such advanced LLMs on-device introduces challenges related to model compression, energy efficiency, and latency. Reinforcement learning (RL) offers a solution by enabling adaptive execution under hardware constraints, dynamically optimizing computational cost and accuracy. Federated reinforcement learning further enhances personalization while maintaining privacy, allowing LLMs to learn from user interactions without cloud dependence. RL can also optimize resource allocation, quantization, and pruning to ensure energy-efficient performance while enabling real-time adaptation to environmental feedback. This makes RL-based LLMs ideal for real-time applications such as autonomous agents and interactive assistants. By integrating RL with On-Device AI, LLMs can become more autonomous, efficient, and privacy-preserving, leading to a new generation of AI systems that seamlessly support ML tasks while operating independently on edge devices.
VII-C Neural-Symbolic LLMs
@@
Current research on LLMs primarily focuses on natural language processing. However, world knowledge extends beyond natural language definitions and expressions. For instance, data in wireless environments often takes non-natural language forms, such as XML and JSON, which are used for data exchange and information structure descriptions to facilitate communication between different systems. In addition, CMOS sensors in mobile phones typically interact with the main control chip through communication interfaces like I2C or SPI, utilizing specific command sets. Therefore, future developments in LLMs should prioritize optimizing models for symbolic languages, enabling them to effectively process and understand various data formats and communication protocols, thereby broadening their applicability across diverse domains. Furthermore, since symbolic languages are implemented through structured rules and representations, they offer clear traceability, logical reasoning capabilities, and enhanced interpretability. Combining symbolic language with LLMs, based on traditional natural language training, can make the reasoning process transparent and easy to understand while accurately expressing complex concepts and relationships.
VII-D Autonomous AI Agents
@@
Autonomous AI agents are intelligent systems built around LLMs as core controllers, enabling them to independently perform tasks, make decisions, and adapt to their environments through reasoning and learning. This vision is realized through models such as AutoGPT [215] and BabyAGI [216], which possess strong general problem-solving capabilities and can function as the brains of autonomous AI agents. However, to fully realize the potential of autonomous AI agents, the evolution from passive models to those capable of reasoning, decision-making, environmental perception, dynamic learning, and continuous updating is essential. For instance, RL techniques can empower these agents to learn from feedback, enhance their responses, and optimize decisions dynamically. In other words, unlike traditional LLMs, RL-enhanced AI agents can plan tasks, solve multi-step problems, and adapt to changing circumstances. By integrating multimodal capabilities, these agents can process text, visual information, and actions, making them suitable for applications such as autonomous research assistants, adaptive tutoring systems, and robotic automation. Furthermore, on-device RL can enhance efficiency by reducing reliance on cloud resources and improving privacy. Future research should focus on developing scalable RL algorithms, hierarchical learning for task planning, and improved exploration strategies to create truly self-learning AI agents that can operate autonomously.
VIII Societal Impact
@@
Ethics and Morality. The advancement of AI and wireless network technologies raises various ethical and moral issues. First, many traditional jobs may be replaced by AI systems that can perform tasks more efficiently and at a lower cost, leading to significant unemployment and exacerbating socioeconomic inequality. Second, unequal access to AI technology may widen the gap in wealth and opportunities. Those who can leverage AI may gain substantial economic benefits, while those without access risk marginalization, resulting in social stratification and unfairness. Third, AI lacks human emotions and empathy, which can lead to decisions misaligned with human values, especially in complex interpersonal and moral situations. Fourth, AI systems may make mistakes and produce illogical or harmful outcomes, potentially undermining the rights of individuals with diminished autonomy due to age or illness. Fifth, AI systems designed to be overly intelligent may create unforeseen and harmful consequences. Sixth, as AI evolves to achieve self-improvement and surpass human intelligence, unpredictable outcomes could emerge, resulting in a loss of human control and significant ethical and social challenges. Finally, as AI becomes more intelligent and autonomous, discussions about whether it should possess rights are increasing, prompting deep reflections on the ethical and legal implications of balancing rights and responsibilities between humans and AI.
@@
Transparency and Accountability. AI systems that leverage the mutual enhancement between wireless networks and AI technologies necessitate a high level of transparency and accountability. Proof-of-concept projects related to AI often stall during testing due to a lack of trust in the results generated by AI models [217]. This distrust arises from several factors, including the opacity of AI decision-making processes and the potential for inherent biases in algorithms. When stakeholders do not clearly understand how decisions are made, they may question the reliability and fairness of AI outputs, leading to hesitant adoption. To address these issues, it is crucial to establish transparent processes that clarify how AI systems operate and make decisions. This can involve providing comprehensive documentation, clear explanations of algorithmic logic, and strategies to identify and mitigate bias. In addition, accountability measures should be implemented to ensure that AI systems comply with ethical standards and are held responsible for their decisions. By fostering transparency and accountability, we can enhance trust in AI technology and facilitate its successful integration into wireless networks and broader societal applications.
@@
Legal Considerations. The integration of AI with wireless networks presents significant legal challenges. When fully AI-driven systems make erroneous or harmful decisions that violate legal standards, determining liability becomes complex, particularly in distributed AI architectures. For instance, when AI systems are developed by multiple organizations or individuals across different regions, the absence of standardized regulations across jurisdictions can lead to inconsistent governance and legal ambiguities. Moreover, privacy and data protection remain critical concerns, as AI systems rely on vast amounts of data, increasing the risk of unauthorized access, misuse, or breaches. The rapid evolution of AI often outpaces legislative efforts, creating regulatory gaps that could be exploited. To mitigate these risks and enable the responsible large-scale deployment of AI systems, it is essential to establish clear legal frameworks, such as robust data protection laws, and continuously adapt regulatory policies in response to AI’s advancing capabilities.
@@
Environmental Sustainability. The rapid advancement of AI and its integration with wireless networks have raised significant environmental concerns, primarily due to the substantial energy consumption required for training and deploying large-scale AI models. For instance, training BigScience’s 176-billion-parameter BLOOM model generated approximately 25 tons of carbon dioxide emissions, which is equivalent to driving a car around the Earth five times [218]. As AI models continue to expand in scale and complexity, their growing carbon footprint poses a serious challenge to environmental sustainability. To mitigate these impacts, it is essential to develop energy-efficient AI architectures, optimize computational processes, and incorporate renewable energy sources into AI training and inference.
IX Conclusion
@@
This survey begins by reviewing recent advancements in both wireless networks and reinforcement learning (RL)-based large language models (LLMs), then takes the first step in exploring how these two domains can mutually empower each other. We provide a comprehensive analysis of this mutual empowerment, including motivations, challenges, and potential solutions. In addition, we examine relevant applications, discuss future research directions, and evaluate the societal impact of this convergence. The primary objective of this survey is to offer readers innovative insights into the interplay between wireless networks and RL-based LLMs, distinguishing our work from existing surveys that primarily focus on LLM applications. Moving forward, we aim to further investigate this topic and provide a more comprehensive analysis.

### v3

@@ 5.4. Personalized Education
Personalized education considers individual learning styles, preferences, and pace to enhance student engagement, foster a deeper interest in learning, and ultimately improve academic performance (Das et al., 2025). The integration of wireless networks and LLMs has the potential to revolutionize personalized education by providing tailored learning experiences for students. Through real-time data transmission from devices, educators can adjust teaching materials based on the progress of different students, addressing diverse learning needs. For example, personalized learning platforms can deliver content suited to various learning styles, with visual learners receiving more graphics and video materials, while auditory learners engage with audio lectures (Zhang, 2024). Teachers can analyze students’ learning data in real-time to identify those needing additional support and provide targeted guidance. In addition, adaptive learning systems can dynamically modify learning paths according to students’ performance, ensuring they learn at an appropriate difficulty level. Furthermore, educational institutions can develop commercial personalized education tools to optimize curriculum settings and resource allocation, promoting the widespread adoption and advancement of personalized education. By leveraging these strategies, the combination of wireless networks and LLM would significantly enhance the effectiveness of personalized education, allowing every student to benefit from a customized learning experience.
6. Future Directions
@@
In this section, we present a vision for the future directions of the development of LLMs for wireless networks, including quantum LLMs, on-device LLMs, neural-symbolic LLMs, and embodied AI agents.
6.1. Quantum LLMs
@@
Quantum computing (Gyongyosi and Imre, 2019) is a new computing paradigm that utilizes the principles of quantum mechanics to manipulate quantum information units for computation. By employing quantum bits (qubits) to process information in parallel, it achieves greater computational power and the capacity to solve more complex problems than traditional computing methods (Gyongyosi and Imre, 2019). The advancement of quantum computing is anticipated to have a profound impact on existing technologies, including LLMs. For instance, quantum computing has the potential to significantly improve the efficiency of computing, improve data security, and speed up model training, thereby greatly boosting the performance of LLMs (Behura and Patra, 2025). However, to fully realize the benefits of quantum computing, it is essential to transform conventional computer algorithms into quantum algorithms specifically designed for quantum architectures. Research on quantum-enhanced LLMs needs to prioritize optimizing model architecture design, model training and inference, and resource allocation to maximize the advantages provided by quantum computing (Behura and Patra, 2025). In addition, exploring quantum-inspired neural networks and quantum machine learning frameworks could unlock new paradigms in language understanding, reasoning, and data processing.
6.2. On-Device LLMs
@@
LLMs must evolve beyond simple response generation to deeply integrate into machine learning (ML) workflows while operating efficiently on edge devices (Zhang et al., 2024c). They should assist with complex ML-specific tasks such as model selection, hyperparameter tuning, feature engineering, and AutoML while leveraging multimodal capabilities to process diverse data types, including text, images, and structured datasets (Zhao et al., 2019). However, deploying such advanced LLMs on-device introduces challenges related to model compression, energy efficiency, and latency. Reinforcement learning (RL) offers a solution by enabling adaptive execution under hardware constraints, dynamically optimizing computational cost and accuracy (Chen et al., 2020). Federated reinforcement learning (Qi et al., 2021) further enhances personalization while maintaining privacy, allowing LLMs to learn from user interactions without cloud dependence. RL can also optimize resource allocation, quantization, and pruning to ensure energy-efficient performance while enabling real-time adaptation to environmental feedback (Aung et al., 2024b). This makes RL-based LLMs ideal for real-time applications such as autonomous agents and interactive assistants. By integrating RL with on-device AI, LLMs can become more autonomous, efficient, and privacy-preserving, leading to a new generation of AI systems that seamlessly support ML tasks while operating independently on edge devices.
6.3. Neural-Symbolic LLMs
@@
Current research on LLMs primarily focuses on natural language processing (Kumar, 2024; Yan et al., 2024; Das et al., 2025). However, world knowledge extends beyond natural language definitions and expressions. For instance, data in wireless environments often takes non-natural language forms, such as XML and JSON, which are used for data exchange and information structure descriptions to facilitate communication between different systems. In addition, CMOS sensors in mobile phones typically interact with the main control chip through communication interfaces like I2C or SPI, utilizing specific command sets (Malviya et al., 2020). Therefore, future developments in LLMs should prioritize optimizing models for symbolic languages, enabling them to effectively process and understand various data formats and communication protocols, thereby broadening their applicability across diverse domains. Furthermore, since symbolic languages are implemented through structured rules and representations, they offer clear traceability, logical reasoning capabilities, and enhanced interpretability (Pulicharla, 2025). Combining symbolic language with LLMs, based on traditional natural language training, can make the reasoning process transparent and easy to understand while accurately expressing complex concepts and relationships.
6.4. Embodied AI Agents
@@
Embodied AI agents are intelligent systems built around LLMs as core controllers, enabling them to independently perform tasks, make decisions, and adapt to their environments through reasoning and learning (Zhang et al., 2024a). This vision is realized through models such as AutoGPT (Yang et al., 2023c) and BabyAGI (Bieger et al., 2015), which possess strong general problem-solving capabilities and can function as the brains of autonomous AI agents. However, to fully realize the potential of autonomous AI agents, the evolution from passive models to those capable of reasoning, decision-making, environmental perception, dynamic learning, and continuous updating is essential. For instance, RL techniques can empower these agents to learn from feedback, enhance their responses, and optimize decisions dynamically (Wang and Zou, 2022). In other words, unlike traditional LLMs, RL-enhanced AI agents can plan tasks, solve multi-step problems, and adapt to changing circumstances. By integrating multimodal capabilities, these agents can process text, visual information, and actions, making them suitable for applications such as autonomous research assistants, adaptive tutoring systems, and robotic automation. Furthermore, on-device RL can enhance efficiency by reducing reliance on cloud resources and improving privacy (Pandey et al., 2020; Tan et al., 2022). Future research should focus on developing scalable RL algorithms, hierarchical learning for task planning, and improved exploration strategies to create truly self-learning AI agents that can operate autonomously.
7. Societal Impact
@@ 7.1. Ethics and Morality
The advancement of AI and wireless network technologies raises various ethical and moral issues. First, many traditional jobs may be replaced by AI systems that can perform tasks more efficiently and at a lower cost, leading to significant unemployment and exacerbating socioeconomic inequality (Kumar, 2024). Second, unequal access to AI technology may widen the gap in wealth and opportunities (Farahani and Ghasemi, 2024). Those who can leverage AI may gain substantial economic benefits, while those without access risk marginalization, resulting in social stratification and unfairness. Third, AI lacks human emotions and empathy, which can lead to decisions misaligned with human values, especially in complex interpersonal and moral situations (Parthasarathy et al., 2024). Fourth, AI systems may make mistakes and produce illogical or harmful outcomes, potentially undermining the rights of individuals with diminished autonomy due to age or illness (Das et al., 2025). Fifth, AI systems designed to be overly intelligent may create unforeseen and harmful consequences (Tocchetti et al., 2022). Sixth, as AI evolves to achieve self-improvement and surpass human intelligence, unpredictable outcomes could emerge, resulting in a loss of human control and significant ethical and social challenges (Parthasarathy et al., 2024). Finally, as AI becomes more intelligent and autonomous, discussions about whether it should possess rights are increasing, prompting deep reflections on the ethical and legal implications of balancing rights and responsibilities between humans and AI (Tocchetti et al., 2022).
7.2. Transparency and Accountability
@@
AI systems that leverage the mutual enhancement between wireless networks and AI technologies necessitate a high level of transparency and accountability (Quttainah et al., 2024). Proof-of-concept projects related to AI often stall during testing due to a lack of trust in the results generated by AI models (Afroogh et al., 2024). This distrust arises from several factors, including the opacity of AI decision-making processes and the potential for inherent biases in algorithms. When stakeholders do not clearly understand how decisions are made, they may question the reliability and fairness of AI outputs, leading to hesitant adoption. To address these issues, it is crucial to establish transparent processes that clarify how AI systems operate and make decisions. This can involve providing comprehensive documentation, clear explanations of algorithmic logic, and strategies to identify and mitigate bias. In addition, accountability measures should be implemented to ensure that AI systems comply with ethical standards and are held responsible for their decisions (Farahani and Ghasemi, 2024). By fostering transparency and accountability, we can enhance trust in AI technology and facilitate its successful integration into wireless networks and broader societal applications.
7.3. Legal Considerations
@@
The integration of AI with wireless networks presents significant legal challenges. When fully AI-driven systems make erroneous or harmful decisions that violate legal standards, determining liability becomes complex, particularly in distributed AI architectures (Das et al., 2025). For instance, when AI systems are developed by multiple organizations or individuals across different regions, the absence of standardized regulations across jurisdictions can lead to inconsistent governance and legal ambiguities (Kumar, 2024). Moreover, privacy and data protection remain critical concerns, as AI systems rely on vast amounts of data, increasing the risk of unauthorized access, misuse, or breaches (Farahani and Ghasemi, 2024). The rapid evolution of AI often outpaces legislative efforts, creating regulatory gaps that could be exploited. To mitigate these risks and enable the responsible large-scale deployment of AI systems, it is essential to establish clear legal frameworks, such as robust data protection laws, and continuously adapt regulatory policies in response to AI’s advancing capabilities.
7.4. Environmental Sustainability
@@
The rapid advancement of AI and its integration with wireless networks have raised significant environmental concerns, primarily due to the substantial energy consumption required for training and deploying large-scale AI models (Ding and Shi, 2024). For instance, training BigScience’s 176-billion-parameter BLOOM model generated approximately 25 tons of carbon dioxide emissions, which is equivalent to driving a car around the Earth five times (Luccioni et al., 2023). As AI models continue to expand in scale and complexity, their growing carbon footprint poses a serious challenge to environmental sustainability. To mitigate these impacts, it is essential to develop energy-efficient AI architectures, optimize computational processes, and incorporate renewable energy sources into AI training and inference.
8. Conclusion
@@
This survey begins by reviewing recent advancements in both wireless networks and RL-based LLMs, then takes the first step in exploring how these two domains can mutually empower each other. We provide a comprehensive analysis of this mutual empowerment, including motivations, challenges, and potential solutions. In addition, we examine relevant applications, discuss future research directions, and evaluate the societal impact of this convergence. The primary objective of this survey is to offer readers innovative insights into the interplay between wireless networks and RL-based LLMs, distinguishing our work from existing surveys that primarily focus on LLM applications. Moving forward, we aim to further investigate this topic and provide a more comprehensive analysis.